OpenStudy (anonymous):

|2x-9|+4 <= 7 (using >= and <= to represent the inequalities) |2x-9|+4 <= 7 -4 -4 |2x-9| <= 3 2x-9 <= 3 -9 -9 2x <= -6 2 2 x <= -3 2x-9 >= -3 -9 -9 2x >= -12 2 2 x >= -6

OpenStudy (ash2326):

Do you know what's mod of x |x| ?

OpenStudy (anonymous):

no. :o

OpenStudy (ash2326):

Ok |x| is defined as \[x\ if\ x\ge 0\] or \[ -x\ if\ x\le 0\] So |+2|=2 and |-2|=-(-2)=2 Do you understand this?

OpenStudy (anonymous):

no, i don't know how to do it like that...

OpenStudy (ash2326):

@0mgl0la My friends will help you with this. I gotta go now. Sorry

OpenStudy (anonymous):

ok. bye. thank you, though. :)

OpenStudy (anonymous):

hiiii 0mgl0la Imma try to help :D

OpenStudy (anonymous):

lol :D ok, go for it.

OpenStudy (anonymous):

ahhh okay this is the whole, what you do t one side you have to do to antoher... comprende give me a sec to see if I can figure it out. I haven't done these in 2 years D:

OpenStudy (anonymous):

Is |2x-9|+4 <= 7 the first question?

OpenStudy (anonymous):

|2x+9|+4\[\le\]7 so, yes.

OpenStudy (anonymous):

\[ |2x-9|+4 \le7 \\ |2x-9| \le 3 \\ -3 \le 2x-9 \le 3 \\ 6=-3+9 \le 2x \le 3+9=12 \\ 3 \le x \le 6 \]

OpenStudy (anonymous):

way to steal meh thunder.